# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ IV.5.1 Monochromatic coherent amplitude
Before the evaluations of the mirror displacements, we consider the coherent state of the incident laser. In the conventional gravitational-wave detectors, the state of the optical beam from the light source is in the single-mode coherent state with the complex amplitude
@@
We note that $\alpha(\omega)$ is real. The corresponding electric field with the amplitude (108) of $\alpha(\omega)$ is the continuous monochromatic carrier field with the frequency $\omega_{0}$. The electric field $\hat{E}_{d}$ is given through Eq. (7) as
@@
The quantum expectation value $I$ of the power of the electric field $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$ is given by
@@
The first term in Eq. (112) is the contribution from the classical carrier field $\alpha(\omega)$, and the second term comes from the vacuum fluctuations of the electric field operator $D_{d}^{\dagger}\hat{E}_{d}(t)D_{d}$. The second term diverges as well-known. We neglect it when we estimate the classical power. However, if we take into account the cut-off in the frequency range due to the time bin and maximal observation time as noted in Sec. II.1, the second term is estimated as a negligible term. To evaluate the averaged classical power $I_{0}$ by neglecting the second term and by the time-average of Eq. (112) as
@@
Then, we denote the factor $N$ in the function $\alpha(\omega)$ is given by
$N=\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}.$ (114)
@@ IV.5.2 Final Input-output relation with mirrors’ motion
Now, we consider the input-output relation (107) with the monochromatic condition (IV.5.1). Furthermore, we consider the sideband picture $\omega=\omega_{0}\pm\Omega$ of the input-output relation under this monochromatic condition (IV.5.1). In this sideband picture, the terms include $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, or $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ appears in the expression of the input-output relation. Here, we neglect these terms because their behavior as
@@
These show that the terms of $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ have the behavior of the rapid oscillation with the frequency $2\omega_{0}$ in the time domain. Since we concentrate only on the behavior of the sideband frequency $\Omega$, these rapid oscillation terms are outside of the frequency range of interest. For these reasons, we ignore the terms $\delta(2\omega_{0}\pm\Omega)$, $D_{d}^{\dagger}\hat{Z}_{dif}(2\omega_{0}\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega)D_{d}$ as an approximation [26]. Applying this approximation, the input-output relation (107) is given by
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{D}$ (117)
@@
$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime}))D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right)D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
@@
To complete the evaluation of this input-output relation (117), we have to specify the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators. These are determined through the Heisenberg equation of motion in the next section.
@@ V Equations for mirrors’ motions and their solutions
The purpose of this section is the evaluation of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators. To discuss this evaluation within quantum mechanics, we first discuss a quantum mechanical model of a forced harmonic oscillator in Sec. V.1. In Sec. V.2, we discuss the equations of mirrors’ motion in the proper reference frame [25] whose center is BS of the interferometer. In Sec. V.3, we evaluate the radiation pressure forces that act on each mirror. In Sec. V.4, we derived the explicit form of the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$ of the mirrors’ displacement operators based on the arguments in Sec. V.1. The results of this section are used to evaluate the input-output relation of the Fabri-Pérot interferometer in Sec. VI.
@@ V.1 Quantum Mechanical model for mirror motions and its solutions
$\hat{H}=\frac{\hat{P}^{2}}{2\mu}+\frac{1}{2}\mu m_{p}^{2}\hat{X}^{2}-\hat{F}(t)\hat{X}$. (118)
The position operator $\hat{X}$ corresponds to the mirror displacement operators $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. $\omega_{p}$ corresponds to the pendulum fundamental frequency of mirrors. The operators $\hat{X}(t)$ and $\hat{P}(t)$ satisfy the canonical commutation relation
@@
From the Hamiltonian (118), Heisenberg’s equations of motion are given by
@@ V.2 Equations of motion for each mirror
Now, we consider the equation of motion for EMs and ITMs to evaluate the Fourier transformations $D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}$, $D_{d}\hat{Z}_{com}(\Omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{diff}(\Omega)D_{d}$ in the input-output relation (117). We assume the masses of ITMs are equal to $m_{ITM}$ and those of EMs are equal to $m_{EM}$. Since we apply the proper reference frame whose center is BS, the displacement of BS vanishes, which means that we have to control the external forces to BS vanish. Furthermore, we only consider the motions of mirrors are restricted so that XITM and XEM move only along the $x$-direction and YITM and YEM move only along the $y$-direction and $x$-direction and $y$-direction is orthogonal. In the proper reference frame [25], the effects of gravitational waves are represented by the components $R_{txtx}$, $R_{tyty}$ of the Riemann curvature. Moreover, we denote the radiation pressure force to mirrors, XITM, YITM, XEM, and YEM from the laser in the interferometer by $\hat{F}_{rpXITM}$, $\hat{F}_{rpYITM}$, $\hat{F}_{rpXEM}$, $\hat{F}_{rpYEM}$, respectively. Then, the equations of motion for $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{XEM}$, and $\hat{X}_{YEM}$, we obtain
@@
As the components of the Riemann tensor, we only consider the plus modes of gravitational waves associated with the $x$- and $y$-directions of the interferometer. In this case, the components of $R_{txtx}$ and $R_{tyty}$ the Riemann tensor are given by
@@
where $z$ is the propagation direction of the laser in the interferometer.
@@
$\hat{Z}_{difITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpXITM}(\omega)-{{\mathscr{F}}}_{rpYITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
@@
$\hat{Z}_{dif}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpx}(\omega)-{{\mathscr{F}}}_{rpy}(\omega)}{2(\omega^{2}-\omega_{p}^{2})}$ (155)
@@
In the input-output relation (117), $\hat{Z}_{dif}(\omega)$, $\hat{Z}_{com}(\omega)$, $\hat{Z}_{comITM}(\omega)$, and $\hat{Z}_{difITM}(\omega)$ do not appear in this expression, but $D_{d}^{\dagger}\hat{Z}_{dif}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{com}(\omega)D_{d}$, $D_{d}^{\dagger}\hat{Z}_{comITM}(\omega)D_{d}$, and $D_{d}^{\dagger}\hat{Z}_{difITM}(\omega)D_{d}$ appear in Eq. (117). Through Eqs. (154) and (155), $\hat{Z}_{dif}(\omega)$ and $\hat{Z}_{com}(\omega)$ are linearly related to ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$. Furthermore, Eqs. (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx) and (LABEL:eq:Eq-of-motion-XYITM-dif-Fourier-approx) yields that $\hat{Z}_{comITM}(\omega)$ and $\hat{Z}_{difITM}(\omega)$ linearly depends on ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$. Therefore, we have to evaluate $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$ to the direct evaluation of the input-output relation (117). These evaluations are discussed in the next subsection.

### v2

@@ IV.5.1 Monochromatic coherent amplitude
Before evaluating the mirror displacements, we consider the coherent state of the incident laser. In the conventional gravitational-wave detectors, the state of the optical beam from the light source is in the single-mode coherent state with the complex amplitude
@@
Here, $N$ is given by
$N=\sqrt{\frac{I_{0}}{\hbar\omega_{0}}},$ (109)
where $I_{0}$ is the averaged classical power of the laser defined by Eq. (114) below. We note that $\alpha(\omega)$ is real. The corresponding electric field with the amplitude (108) of $\alpha(\omega)$ is the continuous monochromatic carrier field with the frequency $\omega_{0}$. The electric field $\hat{E}_{d}$ is given through Eq. (7) as
@@
The quantum expectation value $I$ of the power of the electric field ${{\mathfrak{D}}}_{d}^{\dagger}\hat{E}_{d}(t){{\mathfrak{D}}}_{d}$ is given by
@@
The first term in Eq. (113) is the contribution from the classical carrier field $\alpha(\omega)$, and the second term comes from the vacuum fluctuations of the electric field operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{E}_{d}(t){{\mathfrak{D}}}_{d}$. The second term diverges as well-known. We neglect it when we estimate the classical power. However, if we take into account the cut-off in the frequency range due to the time bin and maximal observation time as noted in Sec. II.1, the second term is estimated as a negligible term. To evaluate the averaged classical power $I_{0}$ by neglecting the second term and by the time-average of Eq. (113) as
@@
Thus, we denote the factor $N$ in the function $\alpha(\omega)$ is given by Eq. (109).
@@ IV.5.2 Final Input-output relation with mirrors’ motion
Now, we consider the input-output relation (107) with the monochromatic condition (IV.5.1). Furthermore, we consider the sideband picture $\omega=\omega_{0}\pm\Omega$ of the input-output relation under this monochromatic condition (IV.5.1). In this sideband picture, the terms include $\delta(2\omega_{0}\pm\Omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$, or ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ appears in the expression of the input-output relation. Here, we neglect these terms because they behave as
@@
These show that the terms of ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ have the behavior of the rapid oscillation with the frequency $2\omega_{0}$ in the time domain. Since we concentrate only on the behavior of the sideband frequency $\Omega$, these rapid oscillation terms are outside of the frequency range of interest. For these reasons, we ignore the terms $\delta(2\omega_{0}\pm\Omega)$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(2\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ as an approximation [31]. Applying this approximation, the input-output relation (107) is given by
${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (117)
@@
$\quad\quad\times\left[i\sin((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos((\omega_{0}\pm\Omega)(\tau_{y}^{\prime}-\tau_{x}^{\prime})){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
@@
$\quad\quad\times\left[i\sin\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\left((2\omega_{0}\pm\Omega)\frac{\tau_{y}^{\prime}-\tau_{x}^{\prime}}{2}\right){{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
@@
To complete the evaluation of this input-output relation (117), we have to specify the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators. These are determined through the Heisenberg equation of motion in the next section.
@@ V Equations for mirrors’ motions and their solutions
The purpose of this section is the evaluation of the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators. To discuss this evaluation within quantum mechanics, we first discuss a quantum mechanical model of a forced harmonic oscillator in Sec. V.1. In Sec. V.2, we discuss the equations of mirrors’ motion in the proper reference frame [30] whose center is BS of the interferometer. In Sec. V.3, we evaluate the radiation pressure forces that act on each mirror. In Sec. V.4, we derived the explicit form of the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$ of the mirrors’ displacement operators based on the arguments in Sec. V.1. The results of this section are used to evaluate the input-output relation of the Fabry-Pérot interferometer in Sec. VI.
@@ V.1 Quantum Mechanical model for mirror motions and its solutions
$\hat{H}=\frac{\hat{P}^{2}}{2\mu}+\frac{1}{2}\mu\omega_{p}^{2}\hat{X}^{2}-\hat{F}(t)\hat{X}$. (118)
The position operator $\hat{X}$ corresponds to the mirror displacement operators $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. $\mu$ and $\omega_{p}$ correspond to the mass and the pendulum fundamental frequency of mirrors, respectively. The operators $\hat{X}(t)$ and $\hat{P}(t)$ satisfy the canonical commutation relation
@@
This assumption (120) is valid for the mirrors’ motion if we ignore the optical spring effects in Appendix B due to the appropriate tuning conditions as discussed in Sec. VIII. From the Hamiltonian (118), Heisenberg’s equations of motion are given by
@@ V.2 Equations of motion for each mirror
Now, we consider the equation of motion for EMs and ITMs to evaluate the Fourier transformations ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ in the input-output relation (117). We assume the masses of ITMs are equal to $m_{ITM}$, and those of EMs are equal to $m_{EM}$. Since we apply the proper reference frame whose center is BS, the displacement of BS vanishes, which means that we have to control the external forces to BS vanish. Furthermore, we only consider the motions of mirrors are restricted so that XITM and XEM move only along the $x$-direction and YITM and YEM move only along the $y$-direction and $x$-direction and $y$-direction is orthogonal. In the proper reference frame, the effects of gravitational waves are represented by the components $R_{txtx}$, $R_{tyty}$ of the Riemann curvature tensor [30]. Moreover, we denote the radiation pressure force to mirrors, XITM, YITM, XEM, and YEM from the laser in the interferometer by $\hat{F}_{rpXITM}$, $\hat{F}_{rpYITM}$, $\hat{F}_{rpXEM}$, $\hat{F}_{rpYEM}$, respectively. Then, the equations of motion for $\hat{X}_{XITM}$, $\hat{X}_{YITM}$, $\hat{X}_{XEM}$, and $\hat{X}_{YEM}$, we obtain
@@
As the components of the Riemann curvature tensor, we only consider the plus modes of gravitational waves associated with the $x$- and $y$-directions of the interferometer. In this case, the components of $R_{txtx}$ and $R_{tyty}$ the Riemann curvature tensor are given by
@@
where $z$ is proper distance to the propagation direction of the laser from BS in the interferometer.
@@
$\hat{Z}_{diffITM}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpXITM}(\omega)-{{\mathscr{F}}}_{rpYITM}(\omega)}{2(\omega^{2}-\omega_{p}^{2})},$
@@
$\hat{Z}_{diff}(\omega)$ $=$ $-\frac{{{\mathscr{F}}}_{rpx}(\omega)-{{\mathscr{F}}}_{rpy}(\omega)}{2(\omega^{2}-\omega_{p}^{2})}$ (155)
@@
In the input-output relation (117), $\hat{Z}_{diff}(\omega)$, $\hat{Z}_{com}(\omega)$, $\hat{Z}_{comITM}(\omega)$, and $\hat{Z}_{diffITM}(\omega)$ do not appear in this expression, but ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\omega){{\mathfrak{D}}}_{d}$ appear in Eq. (117). Through Eqs. (154) and (155), $\hat{Z}_{diff}(\omega)$ and $\hat{Z}_{com}(\omega)$ are linearly related to ${{\mathscr{F}}}_{rpx}(\omega)$ and ${{\mathscr{F}}}_{rpy}(\omega)$. Furthermore, Eqs. (LABEL:eq:Eq-of-motion-XYITM-com-Fourier-approx) and (153) yields that $\hat{Z}_{comITM}(\omega)$ and $\hat{Z}_{diffITM}(\omega)$ linearly depends on ${{\mathscr{F}}}_{rpXITM}(\omega)$ and ${{\mathscr{F}}}_{rpYITM}(\omega)$. Therefore, we have to evaluate ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega){{\mathfrak{D}}}_{d}$ to the direct evaluation of the input-output relation (117). These evaluations are discussed in the next subsection.